OpenStudy (anonymous):

The volume of a certain cylinder, V, varies directly as the SQUARE of the radius, r. If V = 16π when the r = 2, then the equation for the volume, V is V=4 π r2 V = 2 π r2 V = 16 π r2 V=4π over r^2

OpenStudy (lgbasallote):

V varies directly as the square of r so \[V = kr^2\] when v = 16 pi, r = 2 so \[16 \pi = k(2^2)\] \[16 \pi = 4k\] \[\frac{16\pi}{4} = k\] \[4\pi = k\] does that help?

OpenStudy (anonymous):

so is it the fist one

OpenStudy (lgbasallote):

yup